Karen is going to deposits $5000 into an account that pays 6% interest compounded continuously. How much will she have after 5 years?
Answer by rfer(16322)

A=5000(2.7183)^0.06*5
A=5000(2.7183)^0.3
A=5000(1.34986)
A=$6749.30